OpenStudy (anonymous):

Find and simplify (r dot x)(t)? r is -8t^2+40t+128. x is -4t^2+20t+64.

OpenStudy (anonymous):

r = -8t^2+40t+128 x = -4t^2+20t+64 (r.x)(t)=r(x(t))=-8(-4t^2+20t+64)^2+40(-4t^2+20t+64)+128 =-8(16t^4-160t^3-112t^2+2560t+4096) + 40(-4t^2+20t+64) +128 simplify the above and get the answer